OpenStudy (anonymous):

Which of the following equations is the equation of a suitable line of best fit for the data given below? {(45, 18), (62, 13), (72, 11), (55, 15), (38, 21), (26, 31), (93, 9), (64, 13), (29, 28), (33, 25), (49, 17)}

OpenStudy (anonymous):

A.y = 3x + 35 B.y = –10x + 35 C.y=-3/10x-35 D.y=-3/10x+35